OpenStudy (anonymous):

Determine which ordered pair is a solution of y = x + 2/5 (0,0) (-1,2/5) (1,2/5) (2,2 2/5) @texaschic101

pooja195 (pooja195):

Plug the solutions into the equation

pooja195 (pooja195):

(0) = (0) + 2/5 (2/5) = (-1)+ 2/5 (2/5) = (1)+ 2/5 (2 2/5) = (2)+ 2/5 see what makes a true statement

OpenStudy (anonymous):

is it C?

pooja195 (pooja195):

Nope

pooja195 (pooja195):

Try again :)

OpenStudy (anonymous):

I'm....Honestly not sure then...Sorry.

pooja195 (pooja195):

Which ones can we eliminate?

OpenStudy (anonymous):

I know for sure we can eliminate (0, 0)

pooja195 (pooja195):

ok what else?

OpenStudy (anonymous):

B(-1,2/5)

OpenStudy (anonymous):

I know we can eliminate that one

pooja195 (pooja195):

That leaves use with C and D whats 1+ 2/5=?

OpenStudy (anonymous):

1.4

OpenStudy (anonymous):

2 + 2 2/5 = 4.4

pooja195 (pooja195):

good according to you options it equals a difrent number so its incorrect therefore your answer is D (which is what we are left with)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

Thank you!!! :)

pooja195 (pooja195):

:)